OpenStudy (anonymous):

If f(x) = x^2 find f(1*1) - f(1) / (1*1) - (1).

Parth (parthkohli):

\[f(x) = {x^2} \implies {f(1\times1) - f(1) \over (1 \times 1) - 1} \]

Parth (parthkohli):

Is that the question?

OpenStudy (anonymous):

yeah !

Parth (parthkohli):

Look at the denominator, and simplify it.\[ (1 \times 1) - 1 \implies 1 -1=?\]

OpenStudy (anonymous):

okay

Parth (parthkohli):

What is 1 - 1?

OpenStudy (anonymous):

0

Parth (parthkohli):

\[f(1 \times 1) - f(1) \over 0 \]Can you divide by 0?

OpenStudy (anonymous):

no

Parth (parthkohli):

So it is undefined.

OpenStudy (anonymous):

ty :D

Parth (parthkohli):

You're welcome ^_^